# Symmetry Breaking for Inductive Logic Programming

### v2

@@ Abstract
The goal of inductive logic programming is to search for a hypothesis that generalises training data and background knowledge. The challenge is searching vast hypothesis spaces, which is exacerbated because many logically equivalent hypotheses exist. To address this challenge, we introduce a method to break symmetries in the hypothesis space. We implement our idea in answer set programming. Our experiments on multiple domains, including visual reasoning and game playing, show that our approach can reduce solving times from over an hour to just 17 seconds.
@@ 1 Introduction
The fundamental challenge in ILP is searching vast hypothesis spaces. Recent approaches tackle this challenge by delegating the search to an off-the-shelf solver, such as a Boolean satisfiability solver (Ahlgren and Yuen 2013; Hocquette et al. 2024) or an answer set programming (ASP) solver (Law, Russo, and Broda 2014; Kaminski, Eiter, and Inoue 2019; Cropper and Morel 2021; Evans et al. 2021).
@@
In this paper, we address this symmetry challenge. We first define the body-variant problem: determining whether two rules only differ by the naming of body-only variables, such as $r_{1}$ and $r_{2}$ above. We show that this problem is graph-isomorphism hard (Köbler, Schöning, and Torán 1993).
@@
To illustrate our approach, consider rules $r_{1}$ and $r_{2}$ above. Assuming that variables are ordered alphabetically, the literal piece(A,C) of $r_{2}$ has a gap in its arguments because $B$ is missing. Furthermore, size(C,B) contains $B$ but $(A,C)<_{lex}(B,C)$. By renaming $C$ to $B$ and $B$ to $C$ within $r_{2}$ the rule $r_{1}$ results. Observe that, $r_{1}$ contains both $piece(A,B)$ and $size(B,C)$ and $(A,B)<_{lex}(B,C)$. Thus, we can remove $r_{2}$ from the hypothesis space and keep $r_{1}$.
@@
The main novelty of this paper is a method for breaking symmetries in ILP. As far as we are aware, there is no existing work on this topic. We expand on the novelty in Section 2. The impact, which we demonstrate on multiple domains, is vastly reduced solving and learning times. Moreover, as the idea connects multiple AI areas, including machine learning and constraint programming, there is much potential for broad research to advance this idea. Overall, we contribute the following:
@@
  • We experimentally show on multiple domains, including visual reasoning and game playing, that our approach can reduce solving and learning times by 99%.
@@ 2 Related Work
Rule learning. ILP induces rules from data, similar to rule learning methods (Fürnkranz and Kliegr 2015), such as AMIE+ (Galárraga et al. 2015) and RDFRules (Zeman, Kliegr, and Svátek 2021). Rule-mining methods are typically limited to unary and binary relations and require facts as input and operate under an open-world assumption. By contrast, ILP usually operates under a closed-world assumption, supports relations of any arity, and can learn from definite programs as background knowledge.
@@
Constraint-based ILP. Many recent ILP systems frame the ILP problem as a constraint satisfaction problem, such as SAT (Ahlgren and Yuen 2013; Hocquette et al. 2024) and ASP problems (Corapi, Russo, and Lupu 2011; Law, Russo, and Broda 2014; Schüller and Benz 2018; Kaminski, Eiter, and Inoue 2019; Evans et al. 2021; Bembenek, Greenberg, and Chong 2023). Rather than use refinement operators to build a hypothesis literal-by-literal, most of these approaches formulate the ILP problem as a rule selection problem, where they precompute every possible rule in the hypothesis space and then use a solver to search for a subset that generalises the examples. Because they precompute all possible rules, they cannot learn rules with many literals. By contrast, we focus on an approach that does not precompute all rules and instead builds constraints to restrict the generation of rules.
Redundancy in ILP. Symmetry breaking is a form of redundancy elimination. There is much work on avoiding redundancy in ILP, such as when testing hypotheses on training data (Struyf and Blockeel 2003). For instance, Raedt and Ramon (2004) check whether a rule has a redundant atom before testing it on examples to avoid a coverage check. Zeng, Patel, and Page (2014) prune rules with syntactic redundancy. For instance, for the rule h(X) $\leftarrow$ p(X,Y), p(X,Z), the authors detect that p(X,Z) is duplicate to the literal p(X,Y) under the renaming $Z\mapsto Y$, where $Z$ and $Y$ are not in other literals. By contrast, we reduce redundancy through declarative symmetry-breaking constraints.
Symmetry Breaking. Highly symmetric problems, when represented declaratively, can lead solvers, such as ASP and SAT solvers, to perform redundant search. Various ideas for dealing with symmetric search spaces have been proposed. Lex-leader  (Crawford et al. 1996) is a generic approach that uses tools for computing graph symmetries to automatically identify instance-specific symmetries  (McKay 1981; Junttila and Kaski 2007; Darga et al. 2004). Symmetries that are detected are eliminated by introducing lex-leader symmetry-breaking constraints into the declarative specification during a preprocessing phase, before running the solver. This approach has been employed in SAT (Aloul, Markov, and Sakallah 2003; Anders, Brenner, and Rattan 2024) and ASP (Drescher, Tifrea, and Walsh 2011; Devriendt and Bogaerts 2016) as well as other declarative approaches. Tarzariol, Gebser, and Schekotihin (2021) use ILP to learn first-order ASP symmetry-breaking constraints from examples of instance-specific constraints. By contrast, we break symmetries in the ILP hypothesis space.
@@ 3 Inductive Logic Programming
We restrict hypotheses and background knowledge to definite programs with the least Hebrand model semantics (Lloyd 2012). We define a cost function:
@@
Given an ILP input $(E,B,\mathcal{H})$, a cost function $cost_{E,B}~:$$\mathcal{H}~\mapsto$$\mathbb{N}$ assigns a numerical cost to each hypothesis in $\mathcal{H}$.
@@ 4 Complexity of Symmetry Breaking in ILP
Consider the rules in the introduction: $r_{1}$ = zendo(A) $\leftarrow$ piece(A,B), size(B,C), blue(B), small(C) $r_{2}$ = zendo(A) $\leftarrow$ piece(A,C), size(C,B), blue(C), small(B) Observe that $head(r_{1})=head(r_{2})$ and using $\sigma_{1}=\{C\mapsto$ $B,B\mapsto C\}$ and $\sigma_{2}=\{B\mapsto C,C\mapsto B\}$, it follows that $r_{1}\sigma_{1}=r_{2}$ and $r_{2}\sigma_{2}=r_{1}$
@@
Deciding whether two rules are body variants is intractable as there is a reduction from the graph isomorphism problem to the body-variant problem, for which existing decision procedures are computationally prohibitive (Babai and Luks 1983; Babai and Codenotti 2008). The following result illustrates the (GI)-hardness (Köbler, Schöning, and Torán 1993) of the body-variant problem:
@@ 5 Tractable Symmetry Breaking for ILP
Let $l$ be a literal, $args(l_{1})=(x_{1},\dots,x_{n})$, $1\leq i_{1},\cdots,i_{n}\leq n$ such that for all $1\leq j<n$, $x_{i_{j}}\leq_{\mathcal{V}}x_{i_{j+1}}$ and $i_{j}\not=i_{j+1}$. Then $ord\_vars(l)=(x_{i_{1}},\dots,x_{i_{n}})$.
@@
Let $l_{1}=p(D,A,B)$ and $l_{2}=q(C,B)$ be literals where variables are alphabetically ordered and $k=3$. Then $l_{2}<_{lex}^{3}l_{1}$ because $ord\_vars(l_{1})=(A,B,D)$, $ord\_vars(l_{2})=(B,C)$, $\mathit{pre\_pad}_{3}(l_{1})=(A,B,D)$, $\mathit{pre\_pad}_{3}(l_{1})=(x_{*},B,C)$, and $(x_{*},B,C)<_{lex}(A,B,D)$ where $x_{*}=A$.
@@
To simplify our argument we assume that $vars(r)=\{x_{1},\cdots,x_{m}\}$ where for $1\leq j<m$, $x_{j}<_{\mathcal{V}}x_{j+1}$. We prove the proposition by induction, [^3] Induction is performed over a finite total ordering towards a maximum element. selecting the smallest unsafe variable $x$ (according to $_{\mathcal{V}}$) and then constructing a substitution that, when applied to $r$, results in a rule where the smallest unsafe variable is larger than $x$. After finitely many steps the smallest unsafe variable is larger than $x_{m}$, i.e. every variable in the constructed rule is safe. ∎
@@
where the variables are ordered alphabetically. The argument outlined in the proof of Proposition 2 applies to rule $r_{8}$ as follows. Observe that $C$ is the smallest unsafe variable in $r_{8}$. Applying $\sigma_{1}=\{F\mapsto C,C\mapsto F\}\{F\mapsto E\}$ to $r_{8}$ we get the rule:
$r_{9}=h(A,B)\leftarrow p(A,C),p(B,E),p(C,\mathbf{D}).$
Now $D$ is the smallest unsafe variable in $r_{9}$. Applying $\sigma_{2}=\{F\mapsto D,D\mapsto F\}\{F\mapsto E\}$ to $r_{9}$ we get the rule:
$r_{10}=h(A,B)\leftarrow p(A,C),p(B,D),p(C,E).$
@@ 6 Implementation
Although our symmetry-breaking idea is general and system agnostic, we demonstrate it in ASP with the ASP-based ILP system Popper (Cropper and Morel 2021; Cropper and Hocquette 2023a).
@@
There are many versions of Popper, including versions that learn higher-order programs (Purgal, Cerna, and Kaliszyk 2022), learn from noisy data (Hocquette et al. 2024), and learn from probabilistic data (Hillerström and Burghouts 2024). Our symmetry-breaking approach directly works on all these versions. For simplicity, we describe the basic version of Popper.
@@ 7 Experiments
To answer Q1, we compare the solving time (time spent generating hypotheses) of Popper with and without symmetry breaking. We use a solving timeout of 20 minutes per task. To answer Q2, we compare the solving time of Popper with and without symmetry breaking on progressively harder ILP tasks by increasing the number of variables allowed in a rule444 Cropper and Morel (2021) show that the hypothesis space grows exponentially in the number of variables allowed in a rule.. We use a solving timeout of 60 minutes per task. For Q1 and Q2, the independent variable is whether Popper uses symmetry breaking and the dependent variable is the solving time, which depends entirely on the independent variable. To answer Q3, we compare the learning time of Popper with and without symmetry breaking. We use a learning timeout of 60 minutes per task. However, learning times are a function of many things, not only solving time. For instance, symmetry breaking could reduce solving time by 50% allowing Popper to generate twice as many hypotheses but it also needs to test them. Therefore, learning time does not directly measure the impact of symmetry breaking. In other words, the dependent variable (learning time) does not entirely depend on the independent variable.
@@
IGGP. The task is to induce rules from game traces (Cropper, Evans, and Law 2020) from the general game playing competition (Genesereth and Björnsson 2013).
IMDB. A real-world dataset which contains relations about movies (Mihalkova, Huynh, and Mooney 2007).
@@
Figure 2 shows the solving times of Popper with and without symmetry breaking. Significance tests confirm $(p<0.05)$ that symmetry breaking reduces solving times on 97/449 (22%) tasks and increases solving times on 1/449 (0%) tasks. There is no significant difference in the other tasks. The mean decrease in solving time is $178\pm 56$ seconds. The median decrease is 20 seconds with 95% CI between 8 and 51 seconds. The mean and median increase is $41$ seconds. Some improvements are substantial. For instance, for the sokoban-terminal task, symmetry breaking reduces the solving time from $1075\pm 112$ seconds to $59\pm 6$ seconds. These are minimum improvements because Popper without symmetry breaking often times out after 20 minutes. With a longer timeout, we would likely see greater improvements. Overall, the results suggest that the answer to Q1 is yes, symmetry breaking can drastically reduce solving time.
@@
Figure 3 shows the solving times of Popper with and without symmetry breaking on progressively harder ILP tasks. The results show that Popper without symmetry breaking struggles to scale to harder tasks. Without symmetry breaking, the mean solving times when allowed to use 7, 8, or 9 variables in a rule are $125\pm 44$, $1411\pm 310$, and $3600\pm 0$ seconds respectively. By contrast, with symmetry breaking, the mean solving times when allowed to use 7, 8, or 9 variables in a rule are $4\pm 0$, $7\pm 0$, and $17\pm 2$ seconds respectively. In other words, for the hardest task, symmetry breaking reduces solving time from over an hour to only 17 seconds, a 99.5% reduction. Overall, the results suggest that the answer to Q2 is yes, symmetry breaking can reduce solving time when progressively increasing the complexity of tasks.
@@
Figure 4 shows the learning times of Popper with and without symmetry breaking. Significance tests confirm $(p<0.05)$ that symmetry breaking reduces learning times on 128/449 (29%) tasks and increases learning times on 12/449 (3%) tasks. There is no significant difference in the other tasks. The mean decrease in learning time is $385\pm 109$ seconds and the median is 54 seconds with 95% CI between 21 and 134 seconds. The mean increase is $192\pm 121$ seconds and the median is 120 seconds with 95% CI between 54 and 266 seconds. These are minimum improvements because Popper without symmetry breaking often times out after 60 minutes. With a longer timeout, we would likely see greater improvements. Overall, the results suggest that the answer to Q3 is yes, symmetry breaking can drastically reduce learning time.

### v3

@@ Abstract
The goal of inductive logic programming is to search for a hypothesis that generalises training data and background knowledge. The challenge is searching vast hypothesis spaces, which is exacerbated because many logically equivalent hypotheses exist. To address this challenge, we introduce a method to break symmetries in the hypothesis space. We implement our idea in answer set programming. Our experiments on multiple domains, including visual reasoning and game playing, show that our approach can reduce solving times from over an hour to 17 seconds.
Codehttps://github.com/logicand-learning-lab/aaai26-symbreak
@@ 1 Introduction
The fundamental challenge in ILP is searching large hypothesis spaces. Recent approaches tackle this challenge by delegating the search to an off-the-shelf solver, such as a Boolean satisfiability solver (Ahlgren and Yuen 2013; Hocquette et al. 2024) or an answer set programming (ASP) solver (Law et al. 2014; Kaminski et al. 2019; Cropper and Morel 2021; Evans et al. 2021).
@@
In this paper, we address this symmetry challenge. We first define the body-variant problem: determining whether two rules only differ by the naming of body-only variables, such as $r_{1}$ and $r_{2}$ above. We show that this problem is graph-isomorphism hard (Köbler et al. 1993).
@@
To illustrate our approach, consider rules $r_{1}$ and $r_{2}$ above. Assuming that variables are ordered alphabetically, the literal piece(A,C) of $r_{2}$ has a gap in its arguments because $B$ is missing. Furthermore, size(C,B) contains $B$ but $(A,C)<_{lex}(B,C)$. By renaming $C$ to $B$ and $B$ to $C$ within $r_{2}$ the rule $r_{1}$ results. Observe that $r_{1}$ contains both $piece(A,B)$ and $size(B,C)$ and $(A,B)<_{lex}(B,C)$. Thus, we can remove $r_{2}$ from the hypothesis space and keep $r_{1}$.
@@
The main novelty of this paper is a method for breaking symmetries in ILP. As far as we are aware, there is no existing work on this topic. The impact, which we demonstrate on multiple domains, is vastly reduced solving and learning times. Moreover, this work connects multiple AI areas, notably machine learning and constraint programming. Overall, we contribute the following:
@@
  • We show experimentally on multiple domains, including visual reasoning and game playing, that our approach can reduce solving and learning times by 99%.
@@ 2 Related Work
Rule learning. ILP induces rules from data, similar to rule learning methods (Fürnkranz and Kliegr 2015), such as AMIE+ (Galárraga et al. 2015) and RDFRules (Zeman et al. 2021). Rule-mining methods are typically limited to unary and binary relations and require facts as input and operate under an open-world assumption. By contrast, ILP usually operates under a closed-world assumption, supports relations of any arity, and can learn from definite programs as background knowledge.
@@
Constraint-based ILP. Many ILP systems frame the ILP problem as a constraint satisfaction problem, such as SAT (Ahlgren and Yuen 2013; Hocquette et al. 2024) and ASP (Corapi et al. 2011; Law et al. 2014; Schüller and Benz 2018; Kaminski et al. 2019; Evans et al. 2021; Bembenek et al. 2023). Most of these approaches formulate the ILP problem as rule selection where they precompute every possible rule in the hypothesis space and use a solver to search for a subset that generalises the examples. Because they precompute all possible rules, they cannot learn rules with many literals. By contrast, we build constraints to restrict rule generation without precomputation.
Redundancy in ILP. Symmetry breaking is a form of redundancy elimination. There is much work on avoiding redundancy in ILP, such as when testing hypotheses on training data (Struyf and Blockeel 2003). Raedt and Ramon (2004) check whether a rule has a redundant atom before testing it on examples to avoid a coverage check. Zeng et al. (2014) prune rules with syntactic redundancy. For the rule h(X) $\leftarrow$ p(X,Y), p(X,Z), they detect that p(X,Z) duplicates p(X,Y) under the renaming $Z\mapsto Y$, where $Z$ and $Y$ are not in other literals. By contrast, we reduce redundancy through declarative symmetry-breaking constraints.
Symmetry Breaking. Highly symmetric problems, when represented declaratively, can lead solvers such as ASP and SAT to perform redundant search. Lex-leader (Crawford et al. 1996) uses tools for computing graph symmetries (McKay 1981; Junttila and Kaski 2007; Darga et al. 2004) to automatically identify instance-specific symmetries. Detected symmetries are eliminated by introducing lex-leader symmetry-breaking constraints during preprocessing. This approach has been employed in SAT (Aloul et al. 2003; Anders et al. 2024), ASP (Drescher et al. 2011; Devriendt and Bogaerts 2016), and other declarative approaches. Tarzariol et al. (2021) use ILP to learn first-order ASP symmetry-breaking constraints from examples of instance-specific constraints. By contrast, we break symmetries in the ILP hypothesis space. Our notion of symmetry differs from conventional domain-level symmetries (Devriendt et al. 2016). In our work, symmetry refers to syntactic variable interchangeability in rule bodies: two rules are symmetric if one can be obtained from the other by renaming variables.
@@ 3 Inductive Logic Programming
We restrict hypotheses and background knowledge to definite programs with the least Herbrand model semantics (Lloyd 2012). We define a cost function:
@@
Given an ILP input $(E,B,\mathcal{H})$, a cost function $cost_{E,B}:\mathcal{H}\to\mathbb{N}$ assigns a numerical cost to each hypothesis in $\mathcal{H}$.
@@ 4 Complexity of Symmetry Breaking in ILP
Consider the rules in the introduction: $r_{1}$ = zendo(A) $\leftarrow$ piece(A,B), size(B,C), blue(B), small(C) $r_{2}$ = zendo(A) $\leftarrow$ piece(A,C), size(C,B), blue(C), small(B) Observe that $head(r_{1})=head(r_{2})$ and using $\sigma_{1}=\{C\mapsto$ $B,B\mapsto C\}$ and $\sigma_{2}=\{B\mapsto C,C\mapsto B\}$, it follows that $r_{1}\sigma_{1}=r_{2}$ and $r_{2}\sigma_{2}=r_{1}$.
@@
Deciding whether two rules are body variants is intractable as there is a reduction from the graph isomorphism problem to the body-variant problem, for which existing decision procedures are computationally prohibitive (Babai and Luks 1983; Babai and Codenotti 2008). The following result illustrates the (GI)-hardness (Köbler et al. 1993) of the body-variant problem:
@@ 5 Tractable Symmetry Breaking for ILP
Let $l$ be a literal with arguments $args(l)=(x_{1},\dots,x_{n})$. Let $(i_{1},\ldots,i_{n})$ be a permutation of $(1,\ldots,n)$ such that $x_{i_{1}}\leq_{V}x_{i_{2}}\leq_{V}\cdots\leq_{V}x_{i_{n}}$. Then the ordered variables of $l$ are $ord\_vars(l)=(x_{i_{1}},\ldots,x_{i_{n}})$.
@@
Let $l_{1}=p(D,A,B)$ and $l_{2}=q(C,B)$ be literals where variables are alphabetically ordered and $k=3$. Then $l_{2}<_{lex}^{3}l_{1}$ because $ord\_vars(l_{1})=(A,B,D)$, $ord\_vars(l_{2})=(B,C)$, $\mathit{pre\_pad}_{3}(l_{1})=(A,B,D)$, $\mathit{pre\_pad}_{3}(l_{2})=(x_{*},B,C)$, and $(x_{*},B,C)<_{lex}(A,B,D)$ where $x_{*}=A$.
@@
To simplify our argument we assume that $vars(r)=\{x_{1},\cdots,x_{m}\}$ where for $1\leq j<m$, $x_{j}<_{\mathcal{V}}x_{j+1}$. We prove the proposition by induction, selecting the smallest unsafe variable $x$ (according to $_{\mathcal{V}}$) and then constructing a substitution that, when applied to $r$, results in a rule where the smallest unsafe variable is larger than $x$. After finitely many steps the smallest unsafe variable is larger than $x_{m}$, i.e. every variable in the constructed rule is safe. ∎
@@
where the variables are ordered alphabetically. The argument outlined in the proof of Proposition 2 applies to rule $r_{8}$ as follows. Observe that $C$ is the smallest unsafe variable in $r_{8}$. Applying $\sigma_{1}=\{E\mapsto C,C\mapsto F\}\{F\mapsto E\}$ to $r_{8}$ we get the rule:
$r_{9}=h(A,B)\leftarrow p(A,C),p(B,E),p(E,\mathbf{D}).$
Now $D$ is the smallest unsafe variable in $r_{9}$. Applying $\sigma_{2}=\{E\mapsto D,D\mapsto F\}\{F\mapsto E\}$ to $r_{9}$ we get the rule:
$r_{10}=h(A,B)\leftarrow p(A,C),p(B,D),p(D,E).$
@@ 6 Implementation
We demonstrate our idea in ASP with the ILP system Popper (Cropper and Morel 2021; Cropper and Hocquette 2023a).
@@
Our symmetry-breaking approach directly works with all variants of Popperhopper,maxsynth,propper. For simplicity, we describe the basic version of Popper.
@@
Our symmetry-breaking encoding adds at most $O(m\cdot n^{2}\cdot k)$ ground rules to the ASP solver, where $m$ is the number of candidate rules in a hypothesis, $n$ is the number of possible variable tuples, and $k$ is the number of possible variables.
@@ 7 Experiments
To answer Q1, we compare the solving time (time spent generating hypotheses) of Popper with and without symmetry breaking. We use a solving timeout of 20 minutes per task. To answer Q2, we compare the solving time of Popper with and without symmetry breaking on progressively harder ILP tasks by increasing the number of variables allowed in a rule [^3] Cropper and Morel (2021) show that the hypothesis space grows exponentially in the number of variables allowed in a rule.. We use a solving timeout of 60 minutes per task. For Q1 and Q2, the independent variable is whether Popper uses symmetry breaking and the dependent variable is the solving time, which depends entirely on the independent variable. To answer Q3, we compare the learning time of Popper with and without symmetry breaking. We use a learning timeout of 60 minutes per task. However, learning times are a function of many things, not only solving time. For instance, symmetry breaking could reduce solving time by 50% allowing Popper to generate twice as many hypotheses but it also needs to test them. Therefore, learning time does not directly measure the impact of symmetry breaking. In other words, the dependent variable (learning time) does not entirely depend on the independent variable.
@@
IGGP. The task is to induce rules from game traces (Cropper et al. 2020) from the general game playing competition (Genesereth and Björnsson 2013).
IMDB. A real-world dataset which contains relations about movies (Mihalkova et al. 2007).
@@
Figure 2 shows the solving times of Popper with and without symmetry breaking [^4] Detailed per-task results and corresponding improvements are in the appendix.. Significance tests confirm $(p<0.05)$ that symmetry breaking reduces solving times on 97/449 (22%) tasks and increases solving times on 1/449 (0%) tasks. There is no significant difference in the other tasks. The mean decrease in solving time is $178\pm 56$ seconds. The median decrease is 20 seconds with 95% CI between 8 and 51 seconds. The mean and median increase is $41$ seconds. Some improvements are substantial. For instance, for the sokoban-terminal task, symmetry breaking reduces the solving time from $1075\pm 112$ seconds to $59\pm 6$ seconds. These are minimum improvements because Popper without symmetry breaking often times out after 20 minutes. With a longer timeout, we would likely see greater improvements. Overall, the results show that symmetry breaking can drastically reduce solving time.
@@
Figure 3 shows the solving times of Popper with and without symmetry breaking on progressively harder ILP tasks. The results show that Popper without symmetry breaking struggles to scale to harder tasks. Without symmetry breaking, the mean solving times when allowed to use 7, 8, or 9 variables in a rule are $125\pm 44$, $1411\pm 310$, and $3600\pm 0$ seconds respectively. By contrast, with symmetry breaking, the mean solving times when allowed to use 7, 8, or 9 variables in a rule are $4\pm 0$, $7\pm 0$, and $17\pm 2$ seconds respectively. In other words, for the hardest task, symmetry breaking reduces solving time from over an hour to only 17 seconds, a 99.5% reduction. Overall, the results show that symmetry breaking can reduce solving time when progressively increasing the complexity of tasks.
@@
Figure 4 shows the learning times of Popper with and without symmetry breaking. Significance tests confirm $(p<0.05)$ that symmetry breaking reduces learning times on 128/449 (29%) tasks and increases learning times on 12/449 (3%) tasks. There is no significant difference in the other tasks. The mean decrease in learning time is $385\pm 109$ seconds and the median is 54 seconds with 95% CI between 21 and 134 seconds. The mean increase is $192\pm 121$ seconds and the median is 120 seconds with 95% CI between 54 and 266 seconds. These are minimum improvements because Popper without symmetry breaking often times out after 60 minutes. With a longer timeout, we would likely see greater improvements. Overall, the results show that symmetry breaking can drastically reduce learning time.